How long does it take for an object thrown straight upward on the surface of the moon with an initial velocity of 20 m/s, to reach the peak of its path? (acceleration of gravity on the moon is 2 m/s2)

a.

5 s

b.

10 s

c.

20 s

d.

40 s

e.

None of these
Answer #1

using the equation of linear motion we have

v = u + gt

at maximum height v =0

0 = 20 + (-2)*t

t = 20/2 = 10s

so option b) 10 s is the correct answer